# NonlinearIncidence.bib

### v4

@@ Discussion
Nonlinear incidence introduced through having a network structure showed that having network connectivity that was more heterogeneous resulted in a reduction in the upper bound on overshoot and a reduction of the dependence of overshoot on transmission overall. Numerically, the overshoot for a homogeneously connected network (i.e. a complete graph) well-approximates a Kermack-McKendrick ODE model. Thus, an upper bound on the overshoot of approximately $0.3$ in both models [nguyen_fundamental_2023] is perhaps unsurprising. While more heterogeneous networks did not have an overshoot larger than $0.3$, the overshoot was greater at larger $R_{0,network}$ than the homogeneous case. Future studies that incorporate more nuanced aspects of behavior, interventions, and time dynamics in the network model can further elucidate the complexities of epidemics on networks.
@@ Methods
    Choose values of $\sigma$ (heterogeneity), the mean node degree (which can be set through the relationship $\text{Mean}=\lambda\Gamma(1-\frac{1}{\text{log}(\sigma)})$ via $\sigma$ and an appropriate scale parameter ($\lambda$), and $N$ (number of nodes).
@@
    $f(x;\lambda,\sigma)=\frac{-\ln(\sigma)}{\lambda}(\frac{x}{\lambda})^{-\ln(\sigma)-1}e^{-(x/\lambda)^{(}-\ln(\sigma))};x\geq 0;\sigma\in(0,1];\lambda>\mathbb{R}^{+}$
@@ Acknowledgements
Supplemental Materials: Upper Bounds on Overshoot in SIR Models with Nonlinear Incidence (entire supplement moved from main body of text)
@@
Next, let us consider invertible polynomials. Having access to large degree polynomials will allow us to make arbitrarily sharp incidence terms. While the set of invertible polynomials is a relatively small subset of all polynomials, it still contains an infinitely large number of possible functions. Because the procedure will require integration of $f(S)$ while it is in the denominator, the algebraic details of doing this for higher-order polynomials with lower-order terms can quickly become cumbersome. So let us illustrate a test case using just the leading term of a generic cubic function. Let $f(S)=aS^{3}$, where $a\in\mathbb{R}$.
@@
$f(x)=ax^{3}\implies f^{-1}(x)=(\frac{x}{a})^{1/3}$ (26)
$Overshoot$ $=(\frac{2S_{\infty}^{2}}{S_{\infty}+1})^{1/3}-S_{\infty}$ (27)
@@
00 0 $=\frac{1}{3}(\frac{2S_{\infty}^{*2}}{S_{\infty}^{*}+1})^{-2/3}(\frac{(S_{\infty}^{*}+1)\cdot 4S_{\infty}^{*}-2S_{\infty}^{*2}\cdot 1}{(S_{\infty}^{*}+1)^{2}})-1$ (28)
@@
This prediction of the maximal overshoot being independent of $a$, whereas the corresponding critical $R_{0}$ is inversely proportional to $a$ is verified numerical in Figure S1. Furthermore, it can be shown that the independence of the maximal overshoot of $a$ carries holds for any $f(S)=aS^{n}$. Since $a$ can be absorbed into $\beta$, the maximal overshoot is consequently independent of $R_{0}$ and instead only depends on the power of $S$.
Figure S1: The overshoot as a function of $R_{0}$ for an SIR model with nonlinear incidence term of $\beta(aS^{3})I$ for different values of $a$. The dashed horizontal line for $Overshoot^{*}$ and the dashed vertical lines given by $R_{0}^{*}=1/a$ are the theoretical predictions given by the calculations in the text. The solid curves are obtained from numerical simulations using the value of the $a$ parameter.
@@
from which it follows on integration using the substitution $u=\text{csc}(aS)+\text{cot}(aS)$ that $I+S+\frac{ln|\text{csc}(aS)+\text{cot}(aS)|}{R_{0}a}$ is constant along all trajectories.
@@
$I_{0}+S_{0}+\frac{ln|\text{csc}(aS_{0})+\text{cot}(aS_{0})|}{R_{0}a}$ $=I_{\infty}+S_{\infty}+\frac{ln|\text{csc}(aS_{\infty})+\text{cot}(aS_{\infty})|}{R_{0}a}$
@@
$\frac{1}{R_{0}}=(I_{\infty}+S_{\infty}-I_{0}-S_{0})\frac{a}{ln\frac{|\text{csc}(aS_{0})+\text{cot}(aS_{0})|}{|\text{csc}(aS_{\infty})+\text{cot}(aS_{\infty})|}}$ (31)
@@
$\frac{1}{R_{0}}$ $=\frac{a(S_{\infty}-1)}{ln(\frac{|\text{csc}(a)+\text{cot}(a)|}{|\text{csc}(aS_{\infty})+\text{cot}(aS_{\infty})|})}$ (32)
@@
$f(x)=\text{sin}(ax)\implies f^{-1}(x)=\frac{arc\text{sin}(x)}{a}$ (33)
For Step D., substituting the expression for $\frac{1}{R_{0}}$ (32) and $f^{-1}$ (33) into the overshoot equation (10) yields:
$Overshoot$ $=\frac{1}{a}arc\text{sin}(\frac{a(S_{\infty}-1)}{ln(\frac{|\text{csc}(a)+\text{cot}(a)|}{|\text{csc}(aS_{\infty})+\text{cot}(aS_{\infty})|})})-S_{\infty}$ (34)
@@
00 0 $=\frac{1}{a}(\frac{1}{\sqrt{1-(\frac{a(S_{\infty}^{*}-1)}{ln(\frac{|\text{csc}(a)+\text{cot}(a)|}{|\text{csc}(aS_{\infty}^{*})+\text{cot}(aS_{\infty}^{*})|})})^{2}}}\frac{ln(\frac{|\text{csc}(a)+\text{cot}(a)|}{|\text{csc}(aS_{\infty}^{*})+\text{cot}(aS_{\infty}^{*})|})\cdot a-a(S_{\infty}^{*}-1)\cdot\frac{\frac{-|\text{csc}(a)+\text{cot}(a)|\cdot(-a\text{cot}(aS_{\infty}^{*})\text{csc}(aS_{\infty}^{*})-acsc^{2}(aS_{\infty}^{*}))}{(|\text{csc}(aS_{\infty}^{*})+\text{cot}(aS_{\infty}^{*})|)^{2}}}{\frac{|\text{csc}(a)+\text{cot}(a)|}{|\text{csc}(aS_{\infty}^{*})+\text{cot}(aS_{\infty}^{*})|}}}{(ln(\frac{|\text{csc}(a)+\text{cot}(a)|}{|\text{csc}(aS_{\infty}^{*})+\text{cot}(aS_{\infty}^{*})|}))^{2}})-1$ (35)
@@
Now consider $a=\frac{2\pi}{3}$, which produces a non-monotonic $f(S)$ over the unit interval (Figure S2a). Solving (35) for $a=\frac{2\pi}{3}$ yields $S_{\infty}^{*}=0.1163...$ Repeating Steps F. and G. when $a=\frac{2\pi}{3}$ yields:
@@
This leads to the question of what the applicable domain of $a$ is. The larger the value of $a$, the stronger the non-monotonicity of $f(S)$ is (Figure S2a). We can first eliminate values based on the second boundary condition (13) which requires that $f(S=1)>\frac{1}{R_{0}}$. Clearly that condition is violated if $f(S=1)$ is not positive, since $\frac{1}{R_{0}}$ is always positive because $\beta$ and $\gamma$ are both positive-definite. Since $f(S)=\text{sin}(aS)$, then $f(S)$ is negative when $a\in[\pi n,2\pi n],n=\mathbb{N}$. Furthermore, since we have a formula for $\frac{1}{R_{0}}$ using (32), we can set up the inequality explicitly.
@@
$\text{sin}(a\cdot 1)$ $>\frac{a(S_{\infty}-1)}{ln(\frac{|\text{csc}(a)+\text{cot}(a)|}{|\text{csc}(aS_{\infty})+\text{cot}(aS_{\infty})|})}$
These result for the different cases of $f(S)=\text{sin}(aS)$ are shown in Figure S2b.
Figure S2: a) $f(S)=\text{sin}(aS)$ for different values of $a$. b) The overshoot as a function of $R_{0}$ for an SIR model with nonlinear incidence term of $\beta(\text{sin}(aS))I$ for different values of $a$. The dashed horizontal lines for $Overshoot^{*}$ and the dashed vertical lines for $R_{0}^{*}$ are the theoretical predictions given by the calculations in the text. The solid curves are obtained from numerical simulations.
@@
Shown is an example from a network with a very homogeneous network distribution ($\sigma=0.000001$, Figure S4) and a moderately heterogeneous network ($\sigma=0.000001$, Figure S5). The time series for a simulation on both networks in shown in Figure S3. Aside from $\sigma$, the other simulation parameters are held constant (i.e. number of nodes (N=100), transmission probability ($\tau=0.1$), recovery probability ($\rho=0.05$), number of initially infected individuals = 1, and mean degree = 7).
Figure S3: The time series for two network simulations that vary only in their heterogeneity parameter ($\sigma$). Left) The network is
Figure S4: Network representation and corresponding degree distribution for a network with $\sigma=0.000001$ (very homogeneous).
Figure S5: Network representation and corresponding degree distribution for a network with $\sigma=0.3$ (moderate heterogeneity).

### v5

@@ Discussion
Nonlinear incidence introduced through having a network structure showed that having network connectivity that was more heterogeneous resulted in a reduction in the upper bound on overshoot and a reduction of the dependence of overshoot on transmission overall. Numerically, the overshoot for a homogeneously connected network (i.e. a complete graph) well-approximates a Kermack-McKendrick ODE model. The ODE model makes a fundamental assumption of a well-mixed population, and similarly, a network with the highest mixing rate would be a complete graph, which a graph approaches as it becomes more homogeneous and the connectivity increases. Thus, an upper bound on the overshoot of approximately $0.3$ in both models [nguyen_fundamental_2023] is perhaps unsurprising. It is interesting to note though, that due to the stochastic nature of the network simulations, occasionally an epidemic on a very homogeneous contact network will exceed the analytical bound in the ODE framework (see Figure 3, the upper quartiles for $blue$ and $purple$ regions exceed $30\%$). While more heterogeneous networks did not have an overshoot larger than $0.3$, the overshoot was greater at larger $R_{0,network}$ than the homogeneous case. Future studies that incorporate more nuanced aspects of behavior, interventions, and time dynamics in the network model can further elucidate the complexities of epidemics on networks.
@@ Methods
    Choose values of $\sigma$ (heterogeneity), the mean node degree (which can be set through the relationship $\text{Mean}=\lambda\Gamma\left(1-\frac{1}{\text{log}(\sigma)}\right)$ via $\sigma$ and an appropriate scale parameter ($\lambda$), and $N$ (number of nodes).
@@
    $f(x;\lambda,\sigma)=\frac{-\text{\text{ln}}(\sigma)}{\lambda}\left(\frac{x}{\lambda}\right)^{-\text{\text{ln}}(\sigma)-1}e^{-(x/\lambda)^{-\text{\text{ln}}(\sigma)}};x\geq 0;\sigma\in(0,1];\lambda>\mathbb{R}^{+}$
@@ Acknowledgements
Supplemental Materials: Upper Bounds on Overshoot in SIR Models with Nonlinear Incidence
@@
Next, let us consider invertible polynomials. Having access to large degree polynomials will allow us to make arbitrarily sharp incidence terms. This allows one to make the growth rate of the epidemic even sharper than the exponential case if needed. Or in general, having access to the toolbox of all polynomials will allow one to pick an appropriately steep incidence term that best matches the data of the particular outbreak being analyzed. While the set of invertible polynomials is a relatively small subset of all polynomials, it still contains an infinitely large number of possible functions. Because the procedure will require integration of $f(S)$ while it is in the denominator, the algebraic details of doing this for higher-order polynomials with lower-order terms can quickly become cumbersome. So let us illustrate a test case using just the leading term of a generic cubic function. Let $f(S)=aS^{3}$, where $a\in\mathbb{R}$.
@@
$f(x)=ax^{3}\implies f^{-1}(x)=\left(\frac{x}{a}\right)^{1/3}$ (4)
For Step D., we substitute the expression for $\frac{1}{R_{0}}$ (3) and $f^{-1}$ (4) into the overshoot equation ($Overshoot=f^{-1}\left(\frac{1}{R_{0}}\right)-S_{\infty}$).
$Overshoot$ $=\left(\frac{2S_{\infty}^{2}}{S_{\infty}+1}\right)^{1/3}-S_{\infty}$ (5)
@@
00 0 $=\frac{1}{3}\left(\frac{2S_{\infty}^{*2}}{S_{\infty}^{*}+1}\right)^{-2/3}\left(\frac{(S_{\infty}^{*}+1)\cdot 4S_{\infty}^{*}-2S_{\infty}^{*2}\cdot 1}{(S_{\infty}^{*}+1)^{2}}\right)-1$ (6)
@@
This prediction of the maximal overshoot being independent of $a$, whereas the corresponding critical $R_{0}$ is inversely proportional to $a$ is verified numerical in Supplemental Figure 1. Furthermore, it can be shown that the independence of the maximal overshoot of $a$ carries holds for any $f(S)=aS^{n}$. Since $a$ can be absorbed into $\beta$, the maximal overshoot is consequently independent of $R_{0}$ and instead only depends on the power of $S$.
Supplemental Figure 1: The overshoot as a function of $R_{0}$ for an SIR model with nonlinear incidence term of $\beta(aS^{3})I$ for different values of $a$. The dashed horizontal line for $Overshoot^{*}$ and the dashed vertical lines given by $R_{0}^{*}=1/a$ are the theoretical predictions given by the calculations in the text. The solid curves are obtained from numerical simulations using the value of the $a$ parameter.
@@
The last example considered here is a periodic function of $S$, namely $sin(aS)$. This might be helpful in a scenario where one gets periodic fluctuations in the incidence rate. This might be relevant for scenarios that try to capture cyclical effects such as seasonality [grassly_seasonal_2006]. Some models have accounted for seasonality in the incidence through incorporating a periodic function into the transmission parameter [altizer_seasonality_2006, aron_seasonality_1984]. Here we make the assumption the periodicity in the incidence enters through the susceptibles.
@@
from which it follows on integration using the substitution $u=\text{csc}(aS)+\text{cot}(aS)$ that $I+S+\frac{\text{ln}|\text{csc}(aS)+\text{cot}(aS)|}{R_{0}a}$ is constant along all trajectories.
@@
$I_{0}+S_{0}+\frac{\text{ln}|\text{csc}(aS_{0})+\text{cot}(aS_{0})|}{R_{0}a}$ $=I_{\infty}+S_{\infty}+\frac{\text{ln}|\text{csc}(aS_{\infty})+\text{cot}(aS_{\infty})|}{R_{0}a}$
@@
$\frac{1}{R_{0}}=(I_{\infty}+S_{\infty}-I_{0}-S_{0})\frac{a}{\text{ln}\frac{|\text{csc}(aS_{0})+\text{cot}(aS_{0})|}{|\text{csc}(aS_{\infty})+\text{cot}(aS_{\infty})|}}$ (9)
@@
$\frac{1}{R_{0}}$ $=\frac{a(S_{\infty}-1)}{\text{ln}\left(\frac{|\text{csc}(a)+\text{cot}(a)|}{|\text{csc}(aS_{\infty})+\text{cot}(aS_{\infty})|}\right)}$ (10)
@@
$f(x)=\text{sin}(ax)\implies f^{-1}(x)=\frac{\text{arcsin}(x)}{a}$ (11)
For Step D., substituting the expression for $\frac{1}{R_{0}}$ (10) and $f^{-1}$ (11) into the overshoot equation ($Overshoot=f^{-1}\left(\frac{1}{R_{0}}\right)-S_{\infty}$) yields:
$Overshoot$ $=\frac{1}{a}\text{arcsin}\left(\frac{a(S_{\infty}-1)}{\text{ln}\left(\frac{|\text{csc}(a)+\text{cot}(a)|}{|\text{csc}(aS_{\infty})+\text{cot}(aS_{\infty})|}\right)}\right)-S_{\infty}$ (12)
@@
00 0 $=\frac{1}{a}\left(\frac{1}{\sqrt{1-\left(\frac{a(S_{\infty}^{*}-1)}{\text{ln}\left(\frac{|\text{csc}(a)+\text{cot}(a)|}{|\text{csc}(aS_{\infty}^{*})+\text{cot}(aS_{\infty}^{*})|}\right)}\right)^{2}}}\frac{\text{ln}\left(\frac{|\text{csc}(a)+\text{cot}(a)|}{|\text{csc}(aS_{\infty}^{*})+\text{cot}(aS_{\infty}^{*})|}\right)\cdot a-a(S_{\infty}^{*}-1)\cdot\frac{\frac{-|\text{csc}(a)+\text{cot}(a)|\cdot(-a\text{cot}(aS_{\infty}^{*})\text{csc}(aS_{\infty}^{*})-acsc^{2}(aS_{\infty}^{*}))}{(|\text{csc}(aS_{\infty}^{*})+\text{cot}(aS_{\infty}^{*})|)^{2}}}{\frac{|\text{csc}(a)+\text{cot}(a)|}{|\text{csc}(aS_{\infty}^{*})+\text{cot}(aS_{\infty}^{*})|}}}{\left(\text{ln}\left(\frac{|\text{csc}(a)+\text{cot}(a)|}{|\text{csc}(aS_{\infty}^{*})+\text{cot}(aS_{\infty}^{*})|}\right)\right)^{2}}\right)-1$ (13)
@@
Now consider $a=\frac{2\pi}{3}$, which produces a non-monotonic $f(S)$ over the unit interval (Supplemental Figure 2a). Solving (13) for $a=\frac{2\pi}{3}$ yields $S_{\infty}^{*}=0.1163...$ Repeating Steps F. and G. when $a=\frac{2\pi}{3}$ yields:
@@
This leads to the question of what the applicable domain of $a$ is. The larger the value of $a$, the stronger the non-monotonicity of $f(S)$ is (Supplemental Figure 2a). We can first eliminate values based on the condition $R_{0}>\frac{1}{f(S)}$. Clearly that condition is violated if $f(S)$ is not positive, since $R_{0}$ is always positive because $\beta$ and $\gamma$ are both positive-definite. Since $f(S)=\text{sin}(aS)$, then $f(S)$ is negative when $a\in[\pi n,2\pi n],n=\mathbb{N}$. Furthermore, since we have a formula for $\frac{1}{R_{0}}$ using (10), we can set up the inequality explicitly.
@@
$\text{sin}(a\cdot 1)$ $>\frac{a(S_{\infty}-1)}{\text{ln}\left(\frac{|\text{csc}(a)+\text{cot}(a)|}{|\text{csc}(aS_{\infty})+\text{cot}(aS_{\infty})|}\right)}$
These result for the different cases of $f(S)=\text{sin}(aS)$ are shown in Supplemental Figure 2b.
Supplemental Figure 2: The overshoot in a SIR model with periodic incidence. a) $f(S)=\text{sin}(aS)$ for different values of $a$. b) The overshoot as a function of $R_{0}$ for an SIR model with nonlinear incidence term of $\beta(\text{sin}(aS))I$ for different values of $a$. The dashed horizontal lines for $Overshoot^{*}$ and the dashed vertical lines for $R_{0}^{*}$ are the theoretical predictions given by the calculations in the text. The solid curves are obtained from numerical simulations.
@@
Shown is an example from a network with a very homogeneous network distribution ($\sigma=0.000001$, Supplemental Figure 4) and a moderately heterogeneous network ($\sigma=0.000001$, Supplemental Figure 5). The time series for a simulation on both networks in shown in Supplemental Figure 3. Aside from $\sigma$, the other simulation parameters are held constant (i.e. number of nodes (N=100), transmission probability ($\tau=0.1$), recovery probability ($\rho=0.05$), number of initially infected individuals = 1, and $\langle k\rangle$= 7).
Supplemental Figure 3: Time series for two network simulations that vary only in their heterogeneity parameter ($\sigma$). Left) the network degree distribution is parameterized by $\sigma=0.000001$ (very homogeneous). Right) The network degree distribution is parameterized by $\sigma=0.3$ (moderate heterogeneity). For both networks, the other simulation parameters are $N=100,\tau=0.1,\rho=0.05$, number of initially infected individuals = 1, and $\langle k\rangle$ = 7.
Supplemental Figure 4: Network representation and corresponding degree distribution for a network with $\sigma=0.000001$ (very homogeneous).
Supplemental Figure 5: Network representation and corresponding degree distribution for a network with $\sigma=0.3$ (moderate heterogeneity).
Supplemental Figure 6: Number line for the heterogeneity parameter ($\sigma$). Demarcated are common distribution shapes and their relative corresponding $\sigma$ value.